OpenStudy (anonymous):

PLEASE HELP!!!!! On question C its the last problem! PLEASE HELP!!!!!! In may 2003, a gas company had the given rate schedule for natural gas usage in single-family residences. Monty service charge $10.75 per therm service charge 1st 40 therms $0.29801/therm Therms over 40 $0.15906/therm Gas charge $0.3575/therm a) what is the charge for using 0 terms in a month? b) what is the charge for using 200 therms in a month? c) construct a funcion that relates the monthly charge for x therms of gas C= ______________ for 0<=x<=40 _______________ for x>40

OpenStudy (mathstudent55):

a) For 0 therms, the charge is only the monthly service charge. b) For 200 therms, you need to add i) the monthly service charge plus ii) the cost of the first 40 therms ($0.29801/therm * 40 therms) plus iii) the cost of the 160 therms above the first 40 therms ($0.15906/therm * 160 therms) plus iv) the gas charge ($0.3575/therm * 200 therms) c) The function has to take into account the fixed monthly cost plus the cost per therm for the first 40 therms for x therms plus the gas cost per therm for x therms.

OpenStudy (anonymous):

i meant 40 therms for the a

OpenStudy (mathstudent55):

Then it's the fixed monthly charge plus $0.29801 * 40 + $0.3575 * 40

OpenStudy (anonymous):

would it be 2.62? for a

OpenStudy (mathstudent55):

Just 0.29801 * 40 is more than that

OpenStudy (anonymous):

i meant 26.22

OpenStudy (mathstudent55):

Monthly service + 0.29801 * 40 + 0.3575 * 40 = $10.75 + $11.92 + $14.30 You forgot the monthly service charge $10.75

OpenStudy (anonymous):

so it would be 36.97 for a and b would be 100.53 right?

OpenStudy (anonymous):

im not sure how to do c

OpenStudy (mathstudent55):

c) The cost is made up of three components. The first component is the monthly service charge. It is one single charge per month and does not vary with how much gas was used. The first component is simply 10.75. The second component is the "per therm service charge." For 0 - 40 therms per month, that charge is $0.29801 per therm. That means if you use 1 therm, you pay $0.29801 * 1, if you use 2 therms, you pay $0.29801 * 2, if you use x therms, you pay $0.29801 * x = 0.29801x The third component is the gas charge which is $0.3575/therm. This is similar to the second component because it also varies according to how many therms of gas were used. If you use 1 therm, you pay $0.3575 * 1, if you use 2 therms, you pay $0.3575 * 2, if you use x therms, you pay $0.3575 * x = 0.3575x The total cost is the sum of the three components.